### v1

@@ Abstract
Arc-Kayles is a game where two players alternate removing two adjacent vertices until no move is left. Introduced in 1978, its computational complexity is still open. More recently, subtraction games, where the players cannot disconnect the graph while removing vertices, were introduced. In particular, Arc-Kayles admits a non-disconnecting variant that is a subtraction game. We study the computational complexity of subtraction games on graphs, proving that they are PSPACE-complete even on very structured graph classes (split, bipartite of any even girth). We prove that Non-Disconnecting Arc-Kayles can be solved in polynomial-time on unicyclic graphs, clique trees, and subclasses of threshold graphs. We also show that a sufficient condition for a second player-win on Arc-Kayles is equivalent to the graph isomorphism problem.
@@ 1 Introduction
Arc-Kayles is one of the many games introduced by Schaefer in his seminal paper on the computational complexity of games [schaefer1978complexity]. It is a two-player, information-perfect, finite vertex deletion game, in which the players alternate removing two adjacent vertices and all their incident edges from a graph. The game ends when the graph is either empty or reduced to an independent graph, and the first player unable to play loses.
While Schaefer proved the PSPACE-completeness of many games, the complexity of Arc-Kayles surprisingly remains open. Some FPT and XP algorithms have been proposed, and the game has been studied on some specific graph classes[huggan2016polynomial], but the results are quite scarce, the game is still generally open on even subdivided stars with three paths.
@@
Non-disconnecting octal games are also called subtraction games, and have also been studied on graphs [dailly2019connected]: for a non-empty set of integers $S$, the connected subtraction game on $S$, $\operatorname{CSG}(S)$, is played by removing connected subgraphs of order $k$ (where $k\in S$) without disconnecting the graph. The non-disconnecting condition of subtraction games allows for better results, compared to Arc-Kayles, on constrained classes such as trees. Note that Non-Disconnecting Arc-Kayles is $\operatorname{CSG}(\{2\})$.
@@
The aim of this paper is to further the study of the computational complexity of Arc-Kayles and of subtraction games, particularly Non-Disconnecting Arc-Kayles. We present results in PSPACE-completeness for subtraction games, an FPT algorithm for Arc-Kayles and several polynomial-time algorithms for Non-Disconnecting Arc-Kayles on structured graph classes.
[…]
@@
Arc-Kayles: a difficult game. As stated above, while Arc-Kayles was introduced in 1978 [schaefer1978complexity], its complexity remains open to this day. However, there have been some algorithmic results. Determining the winner of Arc-Kayles is FPT when parameterized by the number of rounds [lampis2014computational] and the vertex-cover number, and XP when parameterized by neighborhood diversity [yoshiwatari2022winner, hanaka2023winner]. Arc-Kayles on paths is exactly the octal game $\mathbf{0.07}$, and its Grundy sequence is well-known to have period 34 and a preperiod of 68 [guy1956g]. Arc-Kayles on grid graphs is exactly the game Cram [gardner1974mathematical], and thus there are winning strategies for the first player on even-by-odd grids and for the second player on even-by-even grids (by an easy symmetry argument). The game remains open on odd-by-odd grids; Sprague-Grundy values for grids with at most 30 squares [uiterwijk2018construction] and of size $3\times n$ for $n$ up to 21 and a few other odd-by-odd grids [lemoinecomputation, uiterwijk2019solving] have been computed without any regularity appearing. Finally, the game seems to be hard to solve even on trees: an $\mathcal{O}(2^{\frac{n}{2}})$ enumerative algorithm was proposed [hanaka2023winner]. While the game remains open on subdivided stars with three paths, one of which has size 1 [huggan2016polynomial]—by fixing the length of the second path and having the third one change—the authors conjectured that the sequence of Sprague-Grundy values thus obtained would ultimately be periodic. However, the Sprague-Grundy values of Arc-Kayles are unbounded in the general case [dailly2019generalization].
@@
Subtraction games on graphs. Subtraction games have been extensively studied on heaps of counters, and there are several regularity results on them [albert2019lessons, berlekamp2004winning]. It is important to note that these are different rulesets than NimG or GraphNim, game where Nim-heaps are embedded on the nodes or vertices of the graph [DBLP:journals/tcs/Fukuyama03, DBLP:journals/tcs/Fukuyama03a, DBLP:journals/corr/abs-1101-1507]. Subtraction games on graphs were introduced in [dailly2019connected]. They are of particular interest since they restrict the types of moves available on sparser classes (such as trees), and because they cannot be solved through a simple symmetry strategy started by cutting the graph in two isomorphic subgraphs. Hence, their analysis differs fundamentally from the analysis of vertex deletion games that allow disconnecting.
@@
In [dailly2019connected], the authors proved that, by fixing $G$ and a vertex $u$, the Sprague-Grundy values of the sequence of graphs obtained by appending a path of increasing length to $u$ were ultimately periodic. They also found regularity results for the family $\operatorname{CSG}(\{1,\ldots,N\})$ on subdivided stars with three paths, one of which has size 1, in contrast with Arc-Kayles, which they used as base cases for subdivided stars with specific values of $N$. In [beaudou2018octal], the authors gave polynomial-time algorithms for the game $\operatorname{CSG}(\{1,2\})$ on subdivided stars and bistars.
@@
As for the specific case of Non-Disconnecting Arc-Kayles  [^1] A playable version is available at http://kyleburke.info/DB/combGames/nonDisconnectingArcKayles.html., which is the subtraction game $\operatorname{CSG}(\{2\})$, it is known to be polynomial-time solvable on trees, wheels and grids of height at most 3 [dailly2018criticalite].
@@
Our results and outline of the paper. We further expand algorithmic results on Arc-Kayles and subtraction games on graphs. Our first contribution is a general PSPACE-completeness reduction for most subtraction games:
@@
Both of those results are proved in Section 2, where we also show that deciding whether a symmetry strategy can be applied for Arc-Kayles is as hard as the Graph-Isomorphism problem (GI-hard), even on bipartite graphs.
Finally, in Section 3, we give polynomial-time algorithms for deciding the winner of Non-Disconnecting Arc-Kayles on several structured graph classes: unicyclic graphs, clique trees, and subclasses of threshold graphs.
@@ 2 Hardness results for vertex deletion games
  • For every vertex $v\in V$, create a star with center $v^{\prime}$ and $M-1$ leaves, and add the edge $cv^{\prime}$;
  • For every edge $uv\in E$, create a vertex $e_{uv}$, attach $M$ leaves to it, and add the edges $u^{\prime}e_{uv}$ and $v^{\prime}e_{uv}$.
The reduction is depicted on Figure 1. It is clearly polynomial, since we have $|V^{\prime}|=M|V|+(M+1)|E|+M+2$.
@@
First, note that the only possible moves on $G^{\prime}$ consist of removing a vertex $v^{\prime}$ and its attached leaves. Indeed, since $1\not\in S$, it is impossible to play on an induced star of order more than $M$, except if the graph is reduced to exactly a star of order $M+1$. Hence, it is always impossible to remove $c$ or any $e_{uv}$ for $uv\in E$, leaving the stars centered on vertices $v^{\prime}$ as the only possible moves.
Furthermore, if a vertex $u^{\prime}$ has been removed, then it is impossible to remove any vertex $v^{\prime}$ such that $uv\in E$, since we cannot remove the star centered on $e_{uv}$, as doing so would disconnect the graph.
@@
Note that $G^{\prime}$ has girth 4, since for any edge $uv$ in $G$ there is a 4-cycle $u^{\prime}cv^{\prime}e_{uv}$ in $G^{\prime}$, and that $G^{\prime}$ is bipartite, since any odd cycle in $G$ now contains twice as many vertices since each edge of $G$ has been subdivided, and no cycle containing $c$ can be of odd length. Furthermore, the girth of $G$ can be increased (but will remain even) by subdividing every edge enough times and adding $M+1$ leaves to the newly created vertices: a cycle using $c$ will have length $4+2i$ if we subdivide the edges incident with $c$ $i$ times, afterwards we can subdivide the edges of any more smaller cycles in $G^{\prime}$ to increase the girth. This transformation does not affect the proof, since the newly created vertices can never be removed, thus proving that $G^{\prime}$ can have any even girth. ∎
$u$$v$$w$$u^{\prime}$$v^{\prime}$$w^{\prime}$$e_{uv}$$e_{uw}$$e_{vw}$$c$
@@
We remark that the non-disconnecting variant of Node-Kayles  [^2] A playable version is available at http://kyleburke.info/DB/combGames/nonDisconnectingNodeKayles.html., where the selected vertices and their neighbors must not induce a separating set (i.e., a set of vertices that disconnects the graph), is also PSPACE-complete under this definition.
@@
We reduce from Node-Kayles, and consider both Node-Kayles and Non-Disconnecting Node-Kayles as vertex deletion games. Let $G(V,E)$ be a position of Node-Kayles. We construct a position $G^{\prime}$ of Non-Disconnecting Node-Kayles by replacing, in $V$, every edge $uv$ by the following gadget $X_{uv}$:
  • Subdivide $uv$ twice, let $e_{uv}^{1}$ and $e_{uv}^{2}$ be the two vertices thusly created;
@@
The edge gadget $X_{uv}$ is depicted on Figure 2. Afterwards, connect all the $e_{uv}^{1}$’s and $e_{uv}^{2}$’s for every edge $uv\in E$.
Playing Node-Kayles in $G$ is strictly equivalent to playing Non-Disconnecting Node-Kayles in $G^{\prime}$. Indeed, note that no player can ever select a vertex in a copy of $K_{3,3}$, as doing so would disconnect the two other vertices in its part of the copy from the rest of $G^{\prime}$. Furthermore, selecting $u$ in $G^{\prime}$ prevents both players from selecting $v$ for every $uv\in E$ afterwards, as this would disconnect the copy of $K_{3,3}$ in $X_{uv}$ from the rest of $G^{\prime}$. Finally, it is impossible to select a vertex $e_{uv}^{1}$ or $e_{uv}^{2}$ from any gadget $X_{uv}$, as doing so would disconnect the copy of $K_{3,3}$ in the gadget from the rest of $G^{\prime}$. Hence, every move in $G$ has an equivalent move in $G^{\prime}$ and conversely, so the two game trees, and thus outcomes and Grundy values, are the same. ∎
$u$$v$$e_{uv}^{1}$$e_{uv}^{2}$
Figure 2: The edge gadget $X_{uv}$ of the reduction of Theorem 3
For our third reduction, we need to define Avoid True, a game played on positive (without negations) boolean formulas and an assignment of all variables such that the formula is false. (The variables are initially set to false.) Each turn the current player selects one false variable and switches the assignment to true. Players may not switch variables that cause the overall formula to evaluate to true. Under normal play, the last player to move wins. Avoid True is known to be PSPACE-complete on formulas in disjunctive normal form[schaefer1978complexity]. Furthermore, recall that a split graph is a graph where the vertex set can be partitioned into a clique (so every possible edge within the part exists) and an independent set (so no edge exists within this part).
@@
The proof is obtained by reduction from Avoid True.
[…]
  • For any $1\leq i\leq n$, we add $k$ vertices $v_{i}$ and $v_{i}^{1},\dots,v^{k-1}_{i}$ in $V$ and all the edges $(v_{i},v^{j}_{i})$ for $11\leq j\leq k-$ 1 in $E$.
  • For any $1\leq i,j\leq n$ we add the edge $(v_{i},v_{j})$ in $E$.
  • For any $1\leq j\leq m$, we add a vertex $c_{j}$ in $V$, and if $x_{i}\in C_{j}$, we add the edge $(v_{i},c_{j})$ in $E$.
The obtained graph (an example for $k=2$ is depicted on Figure 3) is a split graph, as the $v_{i}$’s are a clique and all the other vertices are an independent set.
@@
Now we prove that the first (second resp.) player wins in Avoid True if and only if he wins in Non-Disconnecting Arc-Kayles.
Suppose that the first (second resp.) player has a winning strategy $\mathcal{S}$ in Avoid True. We define its strategy in Non-Disconnecting Arc-Kayles as follows:
  • Whenever $\mathcal{S}$ wants to claim a variable $x_{i}$, he plays the set $\{v_{i},v_{i}^{1},\dots,v^{k-1}_{i}\}$
  • If his opponent plays a set $\{v_{i},v^{1}_{i},\dots,v^{k-1}_{i}\}$, he considers that the variable $x_{i}$ has been played against $\mathcal{S}$ in Avoid True.
@@
Let $G=(V,E)$ be a graph such that there exists an involutive automorphism $f$ of $G$ such that for all $e\in E$, we have $e\cap f(e)=\emptyset$. Then $G$ is a $\mathcal{P}$-position in Arc-Kayles.
@@
Unlike Arc-Kayles, in Non-Disconnecting Arc-Kayles, there are graphs with such an $f$ where the symmetry strategy does not give a winner, such as the cycle on six vertices: it has this property, but the outcome is $\mathcal{N}$. Indeed, in Non-Disconnecting Arc-Kayles, the move $f(e)$ can be forbidden as it could disconnect the graph.
@@
Let $G$ be a graph. Verifying that a graph $G$ admits an involutive automorphism $f$ such that, for all $e$, $e\cap f(e)=\emptyset$ is GI-hard, even in bipartite graphs.
Proof.
We provide a reduction from the graph isomorphism problem. Let $G_{1},G_{2}$ be two graphs. If $|V(G_{1})|\neq|V(G_{2})|$, let $G^{\prime}=K_{2}$, two vertices connected by an edge. Otherwise, denote by $n=|V(G_{1})|=|V(G_{2})|$, and let $G_{1}^{\prime},G_{2}^{\prime}$ be the graphs obtained from $G_{1}$ and $G_{2}$ obtained by adding a universal vertex $u_{i}$ and a leaf connected to it $v_{i}$ in $G_{i}$. Let $G^{\prime}=G_{1}^{\prime}\cup G_{2}^{\prime}$ We prove that $G_{1}$ and $G_{2}$ are isomorphic if and only if $G^{\prime}$ admits an edge-disjoint involutive automorphism $f$.
Suppose first that $G^{\prime}$ admits an edge-disjoint involutive automorphism $f$. As $u_{1}$ and $u_{2}$ are the only two vertices of degree $n+1$, they have to be the image of the other by $f$. Then, since $v_{1}$ and $v_{2}$ both are leaves and $f(v_{1})=u_{2}$, we can suppose without loss of generality that $f_{1}(v_{1})=v_{2}$. Now, by the hypothesis, as the automorphism conserves connectivity, $f$ induces an isomorphism between $G^{\prime}_{1}$ and $G^{\prime}_{2}$, and therefore between $G_{1}$ and $G_{2}$.
@@
$f^{\prime}(v)=\left\{\begin{array}[]{ll}u_{2}&\mbox{if }v=u_{1}\\ v_{2}&\mbox{if }v=v_{1}\\ f(v)&\mbox{otherwise.}\end{array}\right.$
@@
Put more simply, deciding whether the second player can win by applying a symmetry strategy (such as on even-even grids) is GI-hard.

### v2

@@ Abstract
Arc-Kayles is a game where two players alternate removing two adjacent vertices until no move is left, the winner being the player who played the last move. Introduced in 1978, its computational complexity is still open. More recently, subtraction games, where the players cannot disconnect the graph while removing vertices, were introduced. In particular, Arc-Kayles admits a non-disconnecting variant that is a subtraction game. We study the computational complexity of subtraction games on graphs, proving that they are PSPACE-complete even on very structured graph classes (split, bipartite of any even girth). We give a quadratic kernel for Non-Disconnecting Arc-Kayles when parameterized by the feedback edge number, as well as polynomial-time algorithms for clique trees and a subclass of threshold graphs. We also show that a sufficient condition for a second player-win on Arc-Kayles is equivalent to the graph isomorphism problem.
@@ 1 Introduction
Arc-Kayles is one of the many games introduced by Schaefer in his seminal paper on the computational complexity of games [schaefer1978complexity]. It is a two-player, information-perfect, finite vertex deletion game, in which the players alternate removing two adjacent vertices and all their incident edges from a graph. The game ends when the graph is either empty or reduced to an independent graph, i.e., a graph containing no edges, and the first player unable to play loses. Since the game is fully deterministic, we consider that both players play perfectly.
While Schaefer proved the PSPACE-completeness of many games, the complexity of Arc-Kayles surprisingly remains open. Some FPT and XP algorithms have been proposed, and the game has been studied on some specific graph classes [huggan2016polynomial], but the results are quite scarce, the game is still generally open even on subdivided stars with three paths.
@@
Non-disconnecting octal games, also called subtraction games, have also been studied on graphs [dailly2019connected]: for a non-empty set of integers $S$, the connected subtraction game on $S$, $\operatorname{CSG}(S)$, is played by removing connected subgraphs of order $k$ (where $k\in S$) without disconnecting the graph. The non-disconnecting condition of subtraction games allows for better results, compared to Arc-Kayles, on constrained classes such as trees. Note that Non-Disconnecting Arc-Kayles (for Non-Disconnecting Arc-Kayles) is $\operatorname{CSG}(\{2\})$.
@@
The aim of this paper is to further the study of the computational complexity of subtraction games, particularly Non-Disconnecting Arc-Kayles. We present several PSPACE-completeness results for subtraction games, a sufficient condition for symmetry strategies for Arc-Kayles, and tractability results for Non-Disconnecting Arc-Kayles: an FPT algorithm parameterized by the feedback edge number, i.e., the minimum size of a set $S$ of edges such that GSG-S contains no cycle. as well as several polynomial-time algorithms on structured graph classes.
Combinatorial Game Theory terminology. Arc-Kayles and subtraction games are impartial games, a subset of combinatorial games which have been extensively studied. Since both players have the same possible moves, impartial games can have two possible outcomes: either the first player has a winning strategy (in which case the game has outcome $\mathcal{N}$) or the second player does (in which case the game has outcome $\mathcal{P}$). A position $G^{\prime}$ that can be reached by a move from a position $G$ is called an option of $G$, the set of options of $G$ is denoted by $\operatorname{opt}(G)$. If every option of $G$ has outcome $\mathcal{N}$, then $G$ has outcome $\mathcal{P}$; conversely, if any option has outcome $\mathcal{P}$, then $G$ has outcome $\mathcal{N}$. A refinement of outcomes are the Sprague-Grundy values [grundy1939mathematics, sprague1935mathematische], denoted by $\mathcal{G}(G)$ for a given position $G$. The value can be computed inductively, with $\mathcal{G}(G)=\operatorname{mex}(\operatorname{opt}(G))$, where $\operatorname{mex}(S)$ is the smallest non-negative integer not in $S$. A position $G$ has outcome $\mathcal{P}$ if and only if $\mathcal{G}(G)=0$. Furthermore, Arc-Kayles and connected subtraction games are vertex deletion games, a family of games where the players alternate removing vertices from a graph following some fixed constraints. For more information and history about combinatorial games, we refer the reader to [albert2019lessons, berlekamp2004winning, conway2000numbers, siegel2013combinatorial].
@@
Arc-Kayles: a difficult game. Introduced in 1978 [schaefer1978complexity], Arc-Kayles’s complexity remains open to this day. However, there have been some algorithmic results. Determining the winner of Arc-Kayles is FPT when parameterized by the number of rounds [lampis2014computational] and the vertex-cover number, and XP when parameterized by neighborhood diversity [hanaka2023winner]. Arc-Kayles on paths is exactly the octal game $\mathbf{0.07}$, and its Grundy sequence is well-known to have period 34 and a preperiod of 68 [guy1956g]. Arc-Kayles on grid graphs is exactly the game Cram [gardner1974mathematical], and thus there are winning strategies for the first player on even-by-odd grids and for the second player on even-by-even grids (by an easy symmetry argument). The game remains open on odd-by-odd grids; Sprague-Grundy values for grids with at most 30 squares [uiterwijk2018construction] and of size $3\times n$ for $n$ up to 21 and a few other odd-by-odd grids [lemoinecomputation, uiterwijk2019solving] have been computed without any regularity appearing. Finally, the game seems to be hard to solve even on trees: an $\mathcal{O}(2^{\frac{n}{2}})$ enumerative algorithm was proposed [hanaka2023winner], later improved to $\mathcal{O}(1.3831^{n})$ [hanaka2024faster]. While the game remains open on subdivided stars with three paths, one of which has size 1 [huggan2016polynomial]—by fixing the length of the second path and having the third one change—the authors conjectured that the sequence of Sprague-Grundy values thus obtained would ultimately be periodic. However, the Sprague-Grundy values of Arc-Kayles are unbounded in the general case [dailly2019generalization].
@@
Subtraction games on graphs. Subtraction games have been extensively studied on heaps of counters, and there are several regularity results on them [albert2019lessons, berlekamp2004winning]. It is important to note that these are different rulesets than NimG or GraphNim, games where Nim-heaps are embedded on the nodes or vertices of the graph [DBLP:journals/tcs/Fukuyama03, DBLP:journals/tcs/Fukuyama03a, DBLP:journals/corr/abs-1101-1507]. Subtraction games on graphs were introduced in [dailly2019connected]. They are of particular interest since they restrict the types of moves available on sparser classes (such as trees), and because they cannot be solved through a simple symmetry strategy started by cutting the graph in two isomorphic subgraphs. Hence, their analysis differs fundamentally from the analysis of vertex deletion games that allow disconnecting.
@@
In [dailly2019connected], the authors proved that, by fixing $G$ and a vertex $u$, the Sprague-Grundy values of the sequence of graphs obtained by appending a path of increasing length to $u$ are ultimately periodic. They also found regularity results for the family $\operatorname{CSG}(\{1,\ldots,N\})$ on subdivided stars with three paths, one of which has size 1, in contrast with Arc-Kayles, which they used as base cases for subdivided stars with specific values of $N$. In [beaudou2018octal], the authors gave polynomial-time algorithms for the game $\operatorname{CSG}(\{1,2\})$ on subdivided stars and bistars.
@@
As for the specific case of Non-Disconnecting Arc-Kayles  [^1] A playable version is available at https://kyleburke.info/DB/combGames/nonDisconnectingArcKayles.html., which is the subtraction game $\operatorname{CSG}(\{2\})$, it is known to be polynomial-time solvable on trees, wheels and grids of height at most 3 [dailly2018criticalite].
Our results and outline of the paper.
Our focus here is the complexity of the problem that takes as input a graph and outputs the outcome of the considered game played on this graph.
@@
We further expand algorithmic results on Arc-Kayles and subtraction games on graphs. Our first contribution is a general PSPACE-completeness reduction for most subtraction games:
@@
Both of those results are proved in Section˜2, where we also show that deciding whether a symmetry strategy can be applied for Arc-Kayles is as hard as the Graph-Isomorphism problem (GI-hard), even on bipartite graphs.
Finally, in Section˜3, we give polynomial-time algorithms for deciding the winner of Non-Disconnecting Arc-Kayles on several structured graph classes. Our main result is on parameterized complexity. Note that the parameters used for Arc-Kayles (vertex cover and neighbourhood diversity) are linked to the number of rounds and do not allow for connectivity conditions, and as such would not be a good fit. On the contrary, since Non-Disconnecting Arc-Kayles is trivial on trees, the feedback edge number is a natural parameter:
Theorem 3.
Non-Disconnecting Arc-Kayles has a quadratic kernel parameterized by the feedback edge number of the graph.
We also study clique trees and subclasses of threshold graphs.
@@ 2 Hardness results for vertex deletion games
  • For every vertex $v\in V$, create a star with center $v^{\prime}$ and M1M-1 leaves, and add the edge $(cv^{\prime})$;
  • For every edge $(uv)\in E$, create a vertex $e_{uv}$, attach $M$ leaves to it, and add the edges $(u^{\prime}e_{uv})$ and $(v^{\prime}e_{uv})$.
The reduction is depicted in Figure˜1. It is clearly polynomial, since we have $|V^{\prime}|=M|V|+(M+1)|E|+M+2$.
@@
First, note that the only possible moves on $G^{\prime}$ consist of removing a vertex $v^{\prime}$ and its attached leaves. Indeed, since $1\not\in S$, it is impossible to play on an induced star of order more than $M$, except if the graph is reduced to exactly a star of order $M+1$. Hence, it is always impossible to remove $c$ or any $e_{uv}$ for $(uv)\in E$, leaving the stars centered on vertices $v^{\prime}$ as the only possible moves.
Furthermore, if a vertex $u^{\prime}$ has been removed, then it is impossible to remove any vertex $v^{\prime}$ such that $(uv)\in E$, since we cannot remove the star centered on $e_{uv}$, as doing so would disconnect the graph.
@@
Note that $G^{\prime}$ has girth 4, since for any edge $(uv)$ in $G$ there is a 4-cycle $u^{\prime}cv^{\prime}e_{uv}$ in $G^{\prime}$, and that $G^{\prime}$ is bipartite, since any odd cycle in $G$ now contains twice as many vertices since each edge of $G$ has been subdivided, and no cycle containing $c$ can be of odd length. Furthermore, the girth of $G$ can be increased (but will remain even) by subdividing every edge enough times and adding $M+1$ leaves to the newly created vertices: a cycle using $c$ will have length $4+2i$ if we subdivide the edges incident with $c$ $i$ times, afterwards we can subdivide the edges of any more smaller cycles in $G^{\prime}$ to increase the girth. This transformation does not affect the proof, since the newly created vertices can never be removed, thus proving that $G^{\prime}$ can have any even girth. ∎
$uu^{\prime}vv^{\prime}ww^{\prime}$$e_{uv}$$e_{uw}$$e_{vw}$$c$
@@
We remark that the non-disconnecting variant of Node-Kayles  [^2] A playable version is available at http://kyleburke.info/DB/combGames/nonDisconnectingNodeKayles.html., where the selected vertices and their neighbors must not induce a separating set (i.e., a set of vertices that disconnects the graph), is also PSPACE-complete:
@@
We reduce from Node-Kayles, and consider both Node-Kayles and Non-Disconnecting Node-Kayles as vertex deletion games. Let $G(V,E)$ be a position of Node-Kayles. We construct a position $G^{\prime}$ of Non-Disconnecting Node-Kayles by replacing, in $V$, every edge $(uv)$ by the following gadget $X_{uv}$:
  • Subdivide $(uv)$ twice, let $e_{uv}^{1}$ and $e_{uv}^{2}$ be the two vertices thusly created;
@@
The edge gadget $X_{uv}$ is depicted in Figure˜2. Afterwards, connect all the $e_{uv}^{1}$’s and $e_{uv}^{2}$’s for every edge $(uv)\in E$.
Playing Node-Kayles in $G$ is strictly equivalent to playing Non-Disconnecting Node-Kayles in $G^{\prime}$. Indeed, note that no player can ever select a vertex in a copy of $K_{3,3}$, as doing so would disconnect the two other vertices in its part of the copy from the rest of $G^{\prime}$. Furthermore, selecting $u$ in $G^{\prime}$ prevents both players from selecting $v$ for every $(uv)\in E$ afterwards, as this would disconnect the copy of $K_{3,3}$ in $X_{uv}$ from the rest of $G^{\prime}$. Finally, it is impossible to select a vertex $e_{uv}^{1}$ or $e_{uv}^{2}$ from any gadget $X_{uv}$, as doing so would disconnect the copy of $K_{3,3}$ in the gadget from the rest of $G^{\prime}$. Hence, every move in $G$ has an equivalent move in $G^{\prime}$ and conversely, so the two game trees, and thus outcomes and Grundy values, are the same. ∎
[…]
Figure 2: The edge gadget $X_{uv}$ of the reduction of Theorem˜4
For the next reduction, we need to define Avoid True, a game played on positive (without negations) boolean formulas and an assignment of all variables such that the formula is false. (The variables are initially set to false.) Each turn the current player selects one false variable and switches the assignment to true. Players may not switch variables that cause the overall formula to evaluate to true. Under normal play, the last player to move wins. Avoid True is known to be PSPACE-complete on formulas in disjunctive normal form [schaefer1978complexity]. Furthermore, recall that a split graph is a graph where the vertex set can be partitioned into a clique (so every possible edge within the part exists) and an independent set (so no edge exists within this part).
@@
The proof is obtained by reduction from Avoid True (see Figure˜3).
Let $\varphi=\underset{i=1}{\overset{m}{\bigvee}}C_{i}$ be an Avoid True formula in disjunctive normal form over the variables $x_{1},\dots,x_{n}$. We construct a graph $G=(V,E)$ as follows:
  • For any $1\leq i\leq n$, we include $k$ vertices $v_{i}$ and $v_{i}^{1},\dots,v^{k-1}_{i}$ in $V$ and all the edges $(v_{i}v^{j}_{i})$ for $1\leq j\leq k-1$ in $E$.
  • For any $1\leq i,j\leq n$ we include the edge $(v_{i}v_{j})$ in $E$.
  • For any $1\leq j\leq m$, we include a vertex $c_{j}$ in $V$, and if $x_{i}\in C_{j}$, we add the edge $(v_{i}c_{j})$ to $E$.
The obtained graph (an example for $k=2$ is depicted in Figure˜3) is a split graph, as the $v_{i}$’s are a clique and all the other vertices are an independent set.
@@
We now prove that the first (second resp.) player wins in Avoid True if and only if they win in Non-Disconnecting Arc-Kayles.
Suppose that the first (second resp.) player has a winning strategy $\mathcal{S}$ in Avoid True. We define their strategy in Non-Disconnecting Arc-Kayles as follows:
  • Whenever $\mathcal{S}$ would take a variable $x_{i}$, they play the set $\{v_{i},v_{i}^{1},\dots,v^{k-1}_{i}\}$
  • If their opponent plays a set $\{v_{i},v^{1}_{i},\dots,v^{k-1}_{i}\}$, they consider that the variable $x_{i}$ has been played against $\mathcal{S}$ in Avoid True.
@@
Let $G=(V,E)$ be a graph such that there exists an edge-disjoint involutive automorphism $f$ of $G$. Then $G$ has outcome $\mathcal{P}$ in Arc-Kayles.
@@
Unlike Arc-Kayles, in Non-Disconnecting Arc-Kayles, there are graphs with such an $f$ where the symmetry strategy does not give a winner, such as the cycle on six vertices: it has this property, but the graph has outcome $\mathcal{N}$. Indeed, in Non-Disconnecting Arc-Kayles, the move $f(e)$ can be forbidden as it could disconnect the graph.
@@
Verifying that a graph $G$ admits an involutive automorphism $f$ such that, for all $e$, $e\cap f(e)=\emptyset$ is GI-hard, even in bipartite graphs.
of Theorem˜7.
We provide a reduction from the graph isomorphism problem. Let $G_{1},G_{2}$ be two graphs. If $|V(G_{1})|\neq|V(G_{2})|$, let $G^{\prime}=K_{2}$, two vertices connected by an edge. Otherwise, denote by $n=|V(G_{1})|=|V(G_{2})|$, and let $G_{1}^{\prime},G_{2}^{\prime}$ be the graphs obtained from $G_{1}$ and $G_{2}$ obtained by adding a universal vertex $u_{i}$ and a leaf connected to it $v_{i}$ in $G_{i}$, and by subdividing all the edges of it once. Let $G^{\prime}=G_{1}^{\prime}\cup G_{2}^{\prime}$. Note that $G^{\prime}$ is bipartite thanks to the subdivision of its edges. We prove that $G_{1}$ and $G_{2}$ are isomorphic if and only if $G^{\prime}$ admits an edge-disjoint involutive automorphism $f$.
Suppose first that $G^{\prime}$ admits an edge-disjoint involutive automorphism $f$. As $u_{1}$ and $u_{2}$ are the only two vertices of degree $n+1$, they have to be the image of the other by $f$. Then, since $v_{1}$ and $v_{2}$ both are leaves connected to $u_{1}$ and $u_{2}$ by paths of length $2$, and since $f(u_{1})=u_{2}$, we can suppose without loss of generality that $f_{1}(v_{1})=v_{2}$.
Now, consider $\phi$ the restriction to $G_{1}$ of $f$. If $w\in G_{1}$, since automorphisms preserve distances, we have that $f(\{w\}$ must be at distance $2$ from $f(u_{1})$. Therefore, since $f(u_{1})=u_{2}$, $f(w)$ is at distance $2$ from $u_{2}$ and therefore corresponds to a vertex in $G_{2}$. This proves that the restriction $G_{1}$ of $f$ goes to $G_{2}$, and thus, since $f$ is injective as an automorphism and since $|V(G_{1})|=|V(G_{2})|$, $\phi$ is an isomorphism between $G_{1}$ and $G_{2}$
@@
[…]
@@
Put more simply, deciding whether the second player can win a game of Arc-Kayles by applying a symmetry strategy is GI-hard in the general case.